What is the electric potential 15.0 cm from a point charge?


1
Expert's answer
2022-03-27T10:56:00-0400

The electric potential "V" at the distance "r" from the point charge "q" is given by

"V=k\frac{q}{r}"

"V=9*10^9\: {\rm N\cdot m^2/C}*\frac{q\: \rm C}{0.15\:\rm m}"